OpenStudy (anonymous):

Could someone help me solve this, please? I don't know how to solve it. 2|3x – 5| – 8 = 8

OpenStudy (anonymous):

Try the equation 9x^2-30x+25 = 144

hero (hero):

2|3x – 5| – 8 = 8 Add 8 to both sides to get 2|3x - 5| = 8 + 8 2|3x - 5| = 16 Divide both sides by two to get: |3x - 5| = 16/2 |3x - 5| = 8 Finish solving from there.

OpenStudy (anonymous):

Thanks so much! xxx

OpenStudy (anonymous):

8|x+3/4|<2 So far I have x<-1/2 But, I'm not sure of the other answer

OpenStudy (anonymous):

\[2|3x – 5| – 8 = 8 \] \[2(3x – 5) – 8 = \pm 8 \]

hero (hero):

@yahoo, don't do things the hard way bro

hero (hero):

8|x+3/4|<2 This is a different problem, right?

OpenStudy (anonymous):

Nevermind I got -1

OpenStudy (anonymous):

And yes, it was.

hero (hero):

All you have to do is divide both sides by 8: |x+3/4|<2/8 Then use property of absolute value to get x+3/4<±2/8

OpenStudy (anonymous):

A furniture maker uses the specification 19.88 ≤ w ≤ 20.12 for the width w in inches of a desk drawer. Write the specification as an absolute value inequality. (1 point) |w – 0.24| ≤ 20.12 |w – 20| ≤ 0.12 |w – 20| ≤ 0.24 |w – 0.12| ≤ 20 Could you help me with this one?

hero (hero):

You have to start with the answer choices, then see which one gets you 19.88 ≤ w ≤ 20.12

hero (hero):

Anytime you have something like |x - a| < b, then you can re-write it as -b < x - a < b

hero (hero):

Then solve from there

hero (hero):

So if you have to solve |x - 4| < 6 |x - 4| < 6 -6 < x - 4 < 6 -6 + 4 < x < 6 + 4 -2 < x < 10